Operations Report – May 9th

Crew 261 Operations Report 09-05-2023 Update #1
SOL: 9
Name of person filing report: Julien Villa-Massone
Non-nominal systems:
· Remote observatory
· 1 Helmet, 1 headset
· Some electric outlets in the kitchen
Notes on non-nominal systems: Failed helmet and headset are set aside and labeled as such.
Electric outlets failed on the sink side of the kitchen. Basic troubleshooting consisting of checking breakers did not solve the problem. A cord extender has been installed to temporarily replace these outlets.
Spirit rover used: No
Hours: (before EVA) 220.8
Beginning charge: (Before EVA) 100
Ending charge: (On return from EVA, before recharging) N/A
Currently charging: Yes
Opportunity rover used: Yes
Hours: 125.6
Beginning charge: 100
Ending charge: 68
Currently charging: Yes
Curiosity rover used: Yes
Hours: 236.4
Beginning charge: 100
Ending charge: 59
Currently charging: Yes
Perseverance rover used: No
Hours: 263.1
Beginning charge: 100
Ending charge: N/A
Currently charging: Yes
General notes on rovers: Drove to Lith Canyon. Nominal operation
Summary of Hab operations: Nominal day
WATER USE: rate of 40 gallons / sol today. Trend since mission start of 33.5 gal/day means we will run out of water 2.2 days prior to mission end. Thanks to water resupply, we will now finish the mission with 70 gallons extra if usage continues as per current trend.
Water (static tank): 268 gallons remaining
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: yes. A leak developed due to tank overflow when Crew Engineer was in long EVA. Mission Support was extremely helpful in addressing with the situation. 3 crew members also helped with the process. Toilet issue is now resolved.
Summary of internet: Nominal
Summary of suits and radios: Nominal
Summary of GreenHab operations: Nominal.
WATER USE: 13 gals total
Heater: Off
Supplemental light: Off
Harvest: 755g of tomatoes (orange+red), thym (10g), mint (12g), sage (10g)
Summary of ScienceDome operations: Nominal.
Dual split: Off
POWER report: Nominal. All MPPTs now functional.
Summary of RAM operations: Nominal
Summary of Tunnels maintenance: No zip-ties added.
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nominal.
Questions, concerns and requests to Mission Support: Thank you for your support today.

Copyright © The Mars Society. All rights reserved. | Main Site